OpenStudy (anonymous):

Write an arithmetic series for which S5 = 10. Help I don't know how to do it can someone explain.

OpenStudy (jdoe0001):

well.... do you know what an arithmetic serie is?

OpenStudy (anonymous):

Yes I just know how I can write one with the given info

OpenStudy (jdoe0001):

\(\bf s_5=10\quad \textit{means, the 5th term is 10, so}\\ \quad \\ s_1\quad ,\quad s_2\quad ,\quad s_3\quad ,\quad s_4\quad ,\quad 10\)

OpenStudy (jdoe0001):

now you just need 4 prior numbers that have a common ratio, an added quantity, that will gradually make up 10

OpenStudy (jdoe0001):

let's try say with 1 and common ratio of 2 so \(\bf 1\quad ,\quad 1+2=3\quad ,\quad 3+2=5\quad ,\quad 5+2=7\quad ,\quad 7+2\ne10\) so that didn't work what do you think what else we can try?

OpenStudy (jdoe0001):

maybe a different starting number? or a different common ratio?

OpenStudy (jdoe0001):

should we start with 1 and us +3, or start with 2 and use + 1 or what else?

OpenStudy (anonymous):

1+3=4, 4+3=7, 7+3=10 Like this?

OpenStudy (jdoe0001):

so you're saying start with 1 and use + 3... let's see that \(\bf 1\quad ,\quad 1+3=4\quad ,\quad 4+3=7\quad ,\quad 7+3=10\quad ,\quad 10+3\ne10\) well, you see, our 4TH term is 10, but we need the 5TH term to be the 10 got a bit close though

OpenStudy (anonymous):

0+2=2, 2+2=4 ,4+2=6 ,6+2=8 ,8+2=10

OpenStudy (jdoe0001):

ohhh so you mean start at 2 and use + 2, let's see \(\bf 2\quad ,\quad 2+2=4\quad ,\quad 4+2=6\quad ,\quad 6+2=8\quad ,\quad 8+2=10\\ \quad \\ 2,4,6,8,10\) :)